OpenStudy (anonymous):

Mary spent a total of $321.80 for a party. She spent $200.32 on food, plus an additional $30.37 for each hour of the party. How long was the party? 5 hours 3 hours 4 hours 6 hours

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (kendricklamar2014):

\[$321.80 - $200.32 = $121.48\]

OpenStudy (kendricklamar2014):

\[$30.37 * 5 = ?\] @Sunnyo

OpenStudy (anonymous):

151.85

OpenStudy (kendricklamar2014):

Yes, So 5 hours is not a solution

OpenStudy (kendricklamar2014):

Now, what is:\[$30.37 * 3 = ?\] @Sunnyo

OpenStudy (anonymous):

91.11

OpenStudy (kendricklamar2014):

Yes, so 3 hours is not a Solution

OpenStudy (kendricklamar2014):

What is:\[$30.37 * 4 = ?\]

OpenStudy (kendricklamar2014):

@Sunnyo

OpenStudy (anonymous):

I had shut off my computer.. :(

OpenStudy (anonymous):

So I lost the work I was doing..

OpenStudy (kendricklamar2014):

Oh sorry :(